OpenStudy (anonymous):

If the volume of a sphere is 36π cubic units, what is the radius? 3 units 4√3 units 9 units

OpenStudy (danjs):

What is the formula for the volume of a sphere?

OpenStudy (anonymous):

thats all i got

OpenStudy (danjs):

nope it is (4/3) * pi *r^3

OpenStudy (anonymous):

which one of these 3 units 4√3 units 9 units

OpenStudy (danjs):

yes, we are going to figure that out

OpenStudy (danjs):

Volume of a sphere = \[\frac{ 4 }{ 3 } * \pi *r * r * r\]

OpenStudy (danjs):

They give you the volume = 36pi \[36 \pi = \frac{ 4 }{ 3 } \pi *r^3\]

OpenStudy (danjs):

\[\frac{ 36 * 3 }{ 4 } = r^3\]

OpenStudy (danjs):

cancelled a pie on both sides, multiplied both sides by 3/4

OpenStudy (danjs):

so r^3 = 27, what is the radius?

OpenStudy (anonymous):

hes really smart @DanJS is right :) XD

OpenStudy (danjs):

got it all?